# Cascade Speculative Drafting for Even Faster LLM Inference

### 21 Dec 2023 v2

@@ 6.2 Experimental Results
Table 2 presents the experimental results. In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 37% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 72%.
Additionally, we emphasize the effectiveness of the MaG algorithm. When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 56% speedup on MMLU and a 24% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible. Thus, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
When comparing CS Drafting with two neural models to CS Drafting with one neural model, the former requires additional memory utilization. Despite the latency reduction, CS Drafting with two neural models can potentially incur more computational cost. However, during our experiments, since the draft models had only 2% of the number of parameters of the target model, the memory cost did not outweigh the reduced latency. Therefore, the CS Drafting algorithm with two neural models still uses less computational resources, as it completes each generation round faster.
Another observation is that despite FLAN-T5-small sometimes outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small. This raises a topic for future study: determining the ideal size of draft models for CS Drafting.
@@ 7.1 Efficienct Methods for Language Model Inference
In the era of large language models, the efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
@@ 7.2 Speculative Decoding
Recent works improved the speed of speculative decoding by experimenting with task-specific knowledge distillation (Zhou et al., 2023), or using approaches similar to speculative decoding to achieve a tradeoff between quality and speed (Leviathan et al., 2023; Kim et al., 2023). However, the reduced quality for speed can be undesirable, and task-specific training might not be applicable to many real-world NLP applications such as chatbots; overall, we found no generic improvement for speculative decoding in the current literature.
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) proposed constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
@@ 8 Conclusion
In our work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. Our experiments demonstrate that CS Drafting achieves up to a 72 percent additional speedup over speculative decoding without any additional training, while maintaining the same generation quality as the target model.
@@ Appendix A Max-Gram Implementation
Appendix B Proof and Analysis of Theorem 4.4
We first prove a lemma that helps the proof of theorem 4.4
Lemma B.1.
$\forall a,b,c,d\in\mathbb{R}_{>0},\gamma\in(0,1)$
$\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}\leq\max(\frac{a}{c},\frac{b}{d})$
@@
Without loss of generality, assume $\frac{a}{c}\geq\frac{b}{d}$,
$\frac{a}{c}\geq\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}$
$\Longleftrightarrow a(\gamma c+(1-\gamma)d)\geq c(\gamma a+(1-\gamma)b)$
$\Longleftrightarrow(1-\gamma)ad\geq(1-\gamma)cb$
$\Longleftrightarrow\frac{a}{c}\geq\frac{b}{d}$
Proof of Theorem 4.4
@@
Let $\alpha_{d_{1}},\alpha_{d2}$ and $c_{d_{1}},c_{d2}$ be the acceptence rate and cost ratio for model $d_{1},d_{2}$ respectively. Without loss of generality, we assume $\alpha_{d_{1}}>\alpha_{d2}$ and $c_{d_{1}}>c_{d2}$ Then the system satisfies $\alpha_{i}=f(i)\alpha_{d_{1}}+(1-f(i))\alpha_{d_{2}}$ and $c_{i}=f(i)\alpha_{d_{1}}+(1-f(i))c_{d_{2}}$ We prove using contradiction. Assuming all optimal systems need at least one interpolation of models at some positions. Let an optimal system of $f$ given $d_{1},d_{2}$, Without loss of generality, assume $r$ to be the first indices such that $f(r)$ not being an integer and the system $f$ to be the one with largest $r$ among all optimal systems.
$T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\alpha_{r}(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+c_{r}+\sum_{i=1,i\neq r}^{k}c_{i}}$
$=\frac{(f(r)\alpha_{d_{1}}+(1-f(r))\alpha_{d_{2}})(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+f(r)c_{d_{1}}+(1-f(r))c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}}$
Let $A_{1}=1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j},A_{2}=\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}$
$=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
$\leq\max{(\frac{\alpha_{d_{1}}A_{1}+A_{2}}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2}}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}})}$
This means that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor that is no less than that of a non-integer $f$, indicating that replacing $f(r)$ with an integer will still yield an optimal system. This contradicts the assumption that $f$ is the system with the largest $r$. As a result, there exists an optimal system that does not require interpolation. ∎
Figure 3: Acceptance rate of draft tokens by their positions on MMLU datasets, evaluated on FLAN-T5-small, base and large models.

### 16 Feb 2024 v3

@@ 5.3 Experimental Results
Comparison to Baselines Table 2 presents the experimental results. In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%. Overall, CS Decoding achieves a 3-4x speedup over autoregressive generation, establishing itself as a faster algorithm over the baselines.
Effectiveness of MaG When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 70% speedup on MMLU and a 32% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible. Notably, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
Draft Model Size Despite FLAN-T5-small mostly outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small. This implies that with the limitation of a single draft model, the ideal size of the draft model might increase with the assistance of the MaG model.
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup $k_{11}$
GSM8K Auto. - 1 -
GSM8K S Decoding 160m $2.48$ 12
GSM8K CS Drafting 160m, mag 2.86 15
MMLU Auto. - 1 -
MMLU S Decoding 160m $2.12$ 10
MMLU CS Drafting 160m, mag 2.64 13
Table 3: The experimental results on LLAMA.
Results of Decoder-only Models As shown in Table 3, the results for decoder-only models exhibit similar traits to those observed in the experiments on encoder-decoder models. Therefore, for speculative decoding-based algorithms, we believe there is no distinct difference between an encoder-decoder model and a decoder-only model, aligning with the results in Leviathan et al. (2023).
@@ 6.1 Efficienct Methods for Language Model Inference
In the era of large language models, efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
@@ 6.2 Speculative Decoding
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) propose constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
@@ 7 Conclusion
In this work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. Our experiments show that CS Drafting achieves up to an 81 percent additional speedup over speculative decoding, while maintaining the same output distribution as the target model. Our work demonstrates that LLM inference can be further sped up by cascades without sacrificing generation quality or requiring additional task-specific training.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Max-Gram Implementation
Appendix B Additonal Results
Figure 3: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on MMLU.
Appendix C Proof
C.1 Proof for Theorem 4.1
@@ Appendix C Proof
The probability of accepting $i$ tokens is $\alpha^{i}-\alpha^{i+1}$, with the exception of the $k+1$-th token, which has a probability of $\alpha^{k}$ of being accepted. This is because it requires all the first $i$ tokens to be accepted and the $i+1$-th token to be rejected for this to happen. Therefore,
$\phi_{(\alpha,k)}(x)=\alpha^{k}x^{k+1}+\sum_{i=0}^{k-1}(\alpha^{i}-\alpha^{i+1})x^{i+1}.$ (3)
By rearranging the terms, we can achieve an expression much easier to work with
$\phi_{(\alpha,k)}(x)=x+\sum_{i=1}^{k}\alpha^{i}(x^{i+1}-x^{i})$ (4)
$=x+(x-1)\sum_{i=1}^{k}\alpha^{i}(x^{i})$ (5)
$=x+(x-1)\frac{1-\alpha^{i+1}x^{i+1}}{1-\alpha x}.$ (6)
C.2 Proof for Theorem 4.3
@@
Let $\alpha^{\prime}=\alpha(\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}})$. We first calculate the expected number of tokens being generated in a step of vertical cascade with $\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}}$. With the property of generating function, the coefficient of term $x^{j}$ of $\phi^{n}(x)$ is the probability of the sum of acceptance length of $n$ speculative step being $j$. Therefore, $\phi^{n}(x)$ represents the probability generating function right before $\mathcal{M}_{t}$ performs the generation.
To achieve the expected number of token acceptances of a probability-generating function, we seek for an operator that can map the probability-generating function into the desired expectation.
To achieve the operator, we begin with a single polynomial term of $x^{j}$. Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ (Leviathan et al., 2023), the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$. In addition, due to the linearity of the operator, this can be extended to any polynomial. Therefore, we achieved the desired operator to map a probability-generating function into the desired expectation.
Apply operator $T_{\alpha}$ to $\phi^{n}(x)$, we achieved the result of $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)}$ for the expected number of accepted tokens. Furthermore, since the number of $\mathcal{M}_{d_{1}}$ calls is $n$, and $\mathcal{M}_{d_{2}}$ is called $k$ time for each $\mathcal{M}_{d_{1}}$ call given a total of $nk$ calls of $\mathcal{M}_{d_{2}}$. The time cost is $1+nc_{d_{1}}+nkc_{d_{2}}$ which implied the EWIF of the system being $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})}$.
For Corollary 4.4, since both $10<\alpha<$ 1 and $10<\alpha^{\prime}<$ 1, we have $\alpha^{i+1}\alpha^{\prime i+1}<\alpha\alpha^{\prime}$, meaning that $\frac{1-\alpha^{k+1}\alpha^{\prime k+1}}{1-\alpha\alpha^{\prime}}<1$. Together with $\alpha-1<0$, we have $\phi_{\alpha^{\prime},k}(\alpha)<1+(\alpha-1)=\alpha$. If we also let $nkc_{d_{2}}=0$, we have $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})}>\frac{1-\alpha\alpha^{n}}{(1-\alpha)(1+nc_{d_{1}})}$, which is the EWIF for speculative decoding with step size $n$.